OpenStudy (loser66):

find equation of the straight line passing through (2,-1), (4,3) I need step to find y-intercept. Please, help

OpenStudy (loser66):

I have y = 1/5 x + b . and don't know how to find b

OpenStudy (phi):

How did you find the slope ? I would do change in y divided by change in x

OpenStudy (phi):

with (2,-1) and (4,3) the y values are -1 and +3 change in y is -1 -3 = -4 the x values are 2 and 4 change in x is 2 -4= -2 change in y divided by change in x is -4/-2 that is not 1/5

OpenStudy (loser66):

yes, y =2x +b.

OpenStudy (phi):

to find b, replace x and y with a point that you know is on the line. you know both (2,-1) and (4,3) are on the line. pick one, and replace x and y in y= 2x+b with the numbers. "solve" for b. can you do that ?

OpenStudy (loser66):

-1 = 2(2) +b ---> b = 5?

OpenStudy (phi):

pretty close. you get -1 = 4 + b if b is 5 you get -1= 4+5 or -1= 9 which is horribly wrong. I would add -4 to both sides to solve for b

OpenStudy (loser66):

-5?

OpenStudy (loser66):

so y = 2x -5 ?

OpenStudy (phi):

yes -1 = 4 + b add -4 to both sides -1 - 4 = 4 - 4 + b -5 = b the answer is y =2x-5 we can check it using the other point, (4,3) if we got it right, (4,3) should work: 3 = 2*4-5 ?? 3 = 8 - 5 3 = 3 yes, that worked.

OpenStudy (loser66):

one more question: how to sketch y =|x+1| -1?

OpenStudy (phi):

you should post it as a new question.

OpenStudy (loser66):

why?

OpenStudy (loser66):

anyway, thanks a lot for the previous one